# arXiv:2602.22038v2[math.AP] 27 Feb 2026

### v2

@@ Abstract
We derive quantitative estimates for the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in the terms of relative entropy. The main novelties lie in the application of the Donsker–Varadhan inequality within the framework of moderately interacting particle systems to handle the nonlinear terms, as well as in the introduction of localization techniques combined with a probabilistic data setting, in order to derive estimates for the quadratic variation terms. Moreover, we prove the existence of a suitable solution to the aforementioned model.
MSC2010 subject classification: 49N90, 60H30, 60K35
@@ 1 Introduction
In this paper, we establish a quantitative derivation of the stochastic 2D vortex model in the whole Euclidean space, which is a particular case of the following stochastic Fokker–Planck equation
@@
where $W_{t}^{i,N}$ and $B_{t}$ are independent standard $\mathbb{R}^{d}$-valued Brownian motions, defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$, $K$ is a singular kernel and $V^{N}$ a suitable scalling.
To the best of the author’s knowledge, this is the first time that (1) is derived for 2D stochastic vortex model, by moderately interacting particles, in the whole space - extending what was known before for bounded kernels in [32], on $\mathbb{R}^{d}$, and in [39], for sub-Coulombian kernels, under periodic boundary conditions - with pathwise quantitative bounds, in the sense of the relative entropy functional.
In the particle system (2), the stochastic term $\sigma_{t}dB_{t}$ accounts for a source of randomness that affects all particles, modeling a shared environmental fluctuation. Models of this nature appear in several contexts, such as problems in statistical mechanics, population dynamics and biology, as well as in the analysis of numerical Monte–Carlo schemes and related applications; see, e.g., [4], [5], [24], [19], and [29].
The main goal of this paper is to investigate the asymptotics in $N$ of the empirical measure defined by
@@
where $\delta_{a}$ is the delta Dirac measure concentrated at $a\in\mathbb{R}^{d}$. For that purpose, considering $V^{N}\doteq N^{\beta}V(N^{\frac{\beta}{d}}\cdot)$, we introduce the mollified empirical measure
@@
in order to apply the relative entropy functional directly to the densities $\rho^{N}$ and $\rho$, since by the Csiszar-Kullback-Pinsker inequality, we get
$\|S^{N}_{t}-\rho_{t}\|_{0}^{2}\lesssim N^{-\frac{\beta}{d}}+\|\rho^{N}_{t}-\rho_{t}\|_{1}^{2}\overset{\text{CKP}}{\lesssim}N^{-\frac{\beta}{d}}+\mathcal{H}\left(\rho^{N}_{t}|\rho_{t}\right),$ (4)
where the metric in the left hand side of (4), is the Kantorovich-Rubinstein metric between two probatility measures. It follows that, if there exists a constant $C=C_{\omega}>0$ such that for any $N\in\mathbb{N}$,
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})\leq C(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}),$ (5)
where $\theta$ is an explicit positive parameter, by (4) and (5), we have, in particular, propagation of chaos for the marginals of the empirical measure of the particle system, see [1, Section 8.3] and [45].
In the present work, we apply for the first time in the context of moderately interacting particles, the Donsker–Varadhan inequality together with the dissipation provided by the Fisher information, to handle the nonlinearity. In addition, we employ localization techniques combined with the probabilistic data setting, to derive estimates for the quadratic variation terms in the whole space, as the estimate established in [39], under periodic boundary conditions, does not apply in the present framework. Moreover, we prove the existence of a suitable solution to the equation (1).
@@
Quantitative estimates for singular interactions have been substantially developed in recent years. The relative entropy method for establishing quantitative propagation of chaos in McKean–Vlasov systems with singular interaction kernels was first introduced in [23]. It was developed in the context of general first-order systems with interaction kernels in $W^{-1,\infty}$, a framework that includes the point vortex model approximating the two-dimensional Navier–Stokes equation. We also refer to [42], where the mean-field limit and quantitative propagation of chaos for McKean–Vlasov equations with singular interaction kernels are established via the modulated energy method. In contrast to the relative entropy approach, which operates at the level of the joint law of the particle system, the modulated energy method is formulated directly in terms of the empirical measure associated with the particles.
@@
Since then, the aforementioned works have paved the way for establishing quantitative propagation of chaos results. In [2] and [3], the strategies given in [23] and [42], were adapted to get quantitative derivation for the Patlak–Keller–Segel model, in optimal subcritical regimes. Additionally, [20], by means of a logarithmic Sobolev inequality, derived uniform in time propagation of chaos for the 2D vortex model, adapting the strategy in [23]. Alternatively, with sharp estimates, global in time derivation of that model, based on relaxation estimates for the limiting equation and inequalities of Riesz modulated energy, was derived by [12], see also [41],[11] and [31], for the modulated energy method. Regarding sub-Coulombian kernels, [6] obtained the derivation of the mean-field approximation for Landau-like equations, while [40] obtained global-in-time mean-field convergence for gradient (conservative) diffusive flows. By extending the BBGKY hierarchy given in [30], for singular kernels, [47] obtained the optimal convergence rate for the relative entropy between the k-particle marginals and the corresponding tensorized law for weakly interacting diffusions with bounded interaction kernels.
@@
With this in view we observe that, the relative entropy method has proven effective for deriving quantitative estimates, for singular kernels, under periodic boundary conditions, while the modulated energy method, which can be applied in the whole space, has primarily been developed for deterministic interaction. However, more recently, quantitative propagation of chaos for the 2D vortex model and the 2D Log Gas on the whole space was obtained in [48] and [49], [7], respectively. In these works, the authors exploit Li–Yau-type estimates and Hamilton-type heat kernel estimates to derive suitable bounds for the limiting solution and its derivatives up to second order, which allows them to extend the strategy given in [23], to $\mathbb{R}^{d}$.
@@
Quantitative estimates in the whole space were derived, in the context of moderately interacting particles  [^2] The moderate interact model was introduced by Oelschlager in [34],[33] and [35]. with singular kernels - including the Biot-Savart, the Keller-Segel and the Dirac delta measure kernels - by [36] and [37], based on a semigroup approach introduced in [17], with further advancements in [8], [21], [38], and [27]. It is important to note that these works do not rely on the relative entropy method, but instead establish quantitative estimates in strong functional topologies. The connection between the time evolution of the relative entropy, of the joint distribution, and moderate interacting regime, on the whole space was carried out by [10], based on a combination between the relative entropy and the regularised $L^{2}$-estimate in [35], deriving a propagation of chaos result for the viscous porous medium equation.
@@
But was in [39], on the torus, that the equation (1), was derived, for singular kernels and common noise, in the context of moderate interaction regime, by applying the relative entropy directly to the mollified empirical measure. Earlier, also the same domain, the results established in [36] were extended to the setting with environmental noise in [26]. It is worth highlighting, that quantitative entropy estimates for systems with individual and common noise, as shown in [43] for incompressible the Navier-Stokes equations, [9] for the Hegselmann-Krause model, and [32] for mean-field systems with bounded kernels. For more results, see [18], [15], [16], and [28].
@@
In order to present our results, this work is organized as follows. In the next two subsections, 1.1 and 1.2, respectively, we introduce the notations and assumptions used throughout the text, setting the stage for the subsequent discussions. In the subsection 1.3, we present our main results. The section 2 is dedicated to the proof of our main results. To ease the presentation, the Appendix, provides the corresponding suitable notion of a solution for the limiting equation, as well as some inequalities used throughout the proofs.
@@ 1.1 Notations
For $d\geq 1$, the space of probability density functions on $\mathbb{R}^{d}$ is denoted by $\mathcal{P}(\mathbb{R}^{d})$. For a normed vector space $U$, we denote its norm by $\|\cdot\|_{U}$, except for $L^{p}$ spaces, where we write $\|\cdot\|_{p}$, for $p\geq 1$. The Kantorovich-Rubinstein metric reads, for any two probability measures $\mu$ and $\nu$ on $\mathbb{R}^{d}$,
@@ 1. $(\mathbf{A}^{V})$
    Let $V^{N}:\mathbb{R}^{d}\to\mathbb{R}$ be the scaling given by $V^{N}(y)=N^{\beta}V(N^{\frac{\beta}{d}}y)$, $\beta\in(0,1)$, with $V:\mathbb{R}^{d}\to\mathbb{R}$ satisfying
@@
    and for $\alpha>d/2$
    $V(y)\leq C_{d}(1+|y|^{2})^{-\alpha}.$ (8)
@@ 4. $(\mathbf{A}^{\rho_{0}})$
    The initial condition $\rho_{0}$ is taken such that $\rho_{0}\in W^{2,\infty}\left(\mathbb{R}^{d}\right)$ and
@@ 5. $(\mathbf{A}^{X_{0}})$
    The initial particle system $(X_{0}^{i,N})_{i=1}^{N}$, $N\in\mathbb{N}$ is a sequence of i.i.d with law $\rho_{0}$.
@@
For an example of mollifier $V:\mathbb{R}^{d}\to\mathbb{R}$ satisfying $\left(\mathbf{A}^{V}\right)$, take
@@
and then, it is divergence in the sense of distributions, of a $L^{\infty}$ matrix $K_{0}$, namely $K=\nabla\cdot K_{0}$, with
$K_{0}(x_{1},x_{2})=\left(\frac{1}{2\pi}\arctan{\left(\frac{x_{1}}{x_{2}}\right)}\right)I.$ (14)
It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$.
@@ 1.3 Statement of the main results
The first result, whose proof is postponed to the Appendix, establishes that the limiting equation is well posed in the appropriate functional framework. Indeed, it provides the regularity properties required in the proof of the main results.
Theorem 1.
Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$. Then, there exist $T_{1}>0$, $C_{3}>0$, such that if $\widetilde{C}_{3}<C_{3}$,
  • there exists a pathwise solution of (1) with initial condition $\rho_{0}$, such that $\rho\in L^{\infty}\left([0,T_{1}];W^{2,1}\cap W^{2,\infty}(\mathbb{R}^{d})\right)$, $\mathbb{P}$-a.s.;
@@
    for all $t\in[0,T_{1}]$ and constants $C_{1},C_{2}$ dependent on $\widetilde{C}_{1}$, $\widetilde{C}_{2}$ and $\widetilde{C}_{3}$.
@@
Assume $(\mathbf{A}^{V})$, $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, $(\mathbf{A}^{\rho_{0}})$, $(\mathbf{A}^{X_{0}})$ and let $\rho$ be a solution of (1), given by Theorem 1. In addition, let the dynamics of the particle system be given by (2) and
$\mathcal{H}(\rho_{0}^{N}|\rho_{0})\leq C,\,\,\,\mathbb{P}-a.s.$
where
@@
Then, there exist a time $T\in(0,T_{1})$ and a constant $C>0$, such that, almost surely for $N>>1$,
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (19)
As a corollary of the previous theorem, we establish propagation of chaos for the marginals of the empirical measure of the particle system. In the sense of [1, Section 8.3], it corresponds to convergence of the empirical measure with respect to the Kantorovich–Rubinstein metric, given in (6).
@@
Under assumptions of Theorem 2, we have
$\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (20)
@@
We observe that, compared with previous works on moderate particle systems without common noise, we lose the order of convergence. This is because the estimation of the martingale term in our work is of order $\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)$, while in [36] is the order $\frac{1}{2}-\frac{\beta}{2}$, see Theorem 1.3 in [36]. When environmental noise is taken into account, the overall framework remains comparable to that of [26]. For the same reason, the range of the parameter $\beta$ is smaller in our work. However, in the present work we obtain an improved convergence rate compared to that established in [39].
2 Proofs of main results
In this section we present the proofs of our main results.
In Subsection 2.1, by applying the Itô’s formula, such as in [39], we derive an evolution equation for the relative entropy functional of the regularized empirical measure with respect to the solution of (1).
In Subsection 2.2, we deal with the quadratic variation terms arising from the application of Itô’s formula.
In Subsection 2.3, we address the nonlinear contributions appearing in the evolution equation for the relative entropy functional obtained in the first step.

### v3

# Quantitative entropy estimates for 2D stochastic vortex model on the whole space under moderate interactions

@@ Abstract
We derive quantitative estimates for the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in terms of relative entropy. The main novelties lie in the application of the Donsker–Varadhan inequality in the framework of moderately interacting particle systems to handle the nonlinear terms, as well as in the introduction of localization techniques combined with the probabilistic data setting, in order to derive estimates for the quadratic variation terms. As an application of our entropy bounds, we derive a new energy estimate for the distance between the regularized empirical measure and the solution to the aforementioned model by combining the control of the Fisher information of the particle system, the Ladyzhenskaya inequality, and a nonlinear Grönwall lemma. Moreover, we prove the existence of a suitable solution to the limiting process.
MSC2010 subject classification: 35R60, 35Q30, 82C22
@@ 1 Introduction
In this work, we derive quantitative entropy estimates for the 2D stochastic vortex model in the whole Euclidean space, which is a particular case of the following stochastic Fokker–Planck equation
@@
where $W_{t}^{i,N}$ and $B_{t}$ are independent standard $\mathbb{R}^{d}$-valued Brownian motions, defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$, $K$ is a singular kernel, $V^{N}$ is a suitable scaling and $\sigma$ is a diffusion matrix.
To achieve these bounds, we study the asymptotic behavior, as $N\to\infty$ of the empirical measure defined by
@@
where $\delta_{a}$ is the Dirac delta measure concentrated at $a\in\mathbb{R}^{d}$, by introducing the mollified empirical measure
@@
which allows us to apply the relative entropy functional directly to the densities $\rho^{N}$ and $\rho$.
To the best of our knowledge, this is the first time that pathwise quantitative bounds, in the sense of the relative entropy functional, have been derived for the 2D stochastic vortex model in the whole space via moderately interacting particles. This extends previously known results for bounded kernels on $\mathbb{R}^{d}$ [33] and for sub-Coulombian kernels under periodic boundary conditions [40]. As a consequence of our entropy bounds, we obtain energy estimates for the difference between the regularized empirical measure and the solution of (1), extending previous results by [27] and [37].
Our main novelty lies in the key technical tools we develop. For the first time in the context of moderately interacting particles, we apply the Donsker–Varadhan inequality, together with the dissipation provided by the Fisher information, to handle the nonlinearity. In addition, we employ localization techniques combined with our probabilistic data setting to derive estimates for the quadratic variation terms in the whole space, since the estimates established in [40] under periodic boundary conditions do not apply in the present framework. Finally, new energy estimates quantifying the distance between the regularized empirical measure and the solution to (1) are derived by combining the control of the Fisher information of the particle system with the Ladyzhenskaya inequality and a nonlinear Grönwall inequality. Moreover, we prove the existence of a suitable solution to the equation (1).
@@
Quantitative estimates for singular interactions have been substantially developed in recent years. The relative entropy method for establishing quantitative propagation of chaos in McKean–Vlasov systems with singular interaction kernels was first introduced in [24]. It was developed in the context of general first-order systems with interaction kernels in $W^{-1,\infty}$, a framework that includes the 2D viscous vortex model. We also refer to [43], where quantitative propagation of chaos for McKean–Vlasov equations with singular interaction kernels are established via the modulated energy method. In contrast to the relative entropy approach, which operates at the level of the joint law of the particle system, the modulated energy method is formulated directly in terms of the empirical measure associated with the particles.
@@
Since then, the aforementioned works have paved the way for establishing quantitative propagation of chaos results. In [2] and [3], the strategies given in [24] and [43], were adapted to obtain a quantitative derivation for the Patlak–Keller–Segel model, in optimal subcritical regimes. Additionally, [21], by means of a logarithmic Sobolev inequality, derived uniform in time propagation of chaos for the 2D vortex model, adapting the strategy in [24]. Alternatively, with sharp estimates, global in time derivation of that model, based on relaxation estimates for the limiting equation and inequalities of Riesz modulated energy, was derived by [12], see also [11], [32] and [42], for the modulated energy method. Regarding sub-Coulombian kernels, [6] obtained the derivation of the mean-field approximation for Landau-like equations, while [41] obtained global-in-time mean-field convergence for gradient (conservative) diffusive flows. By extending the BBGKY hierarchy introduced in [31] to the case of singular kernels, and adapting the strategy developed in [24], [48] obtained the optimal convergence rate for the relative entropy between the k-particle marginals and the corresponding tensorized law.
@@
More recently, quantitative propagation of chaos for the 2D vortex model on the whole space was obtained in [49, 50], while the corresponding result for the 2D log-gas was established in [7]. In these works, the authors exploit Li–Yau-type estimates and Hamilton-type heat kernel estimates to derive suitable bounds for the limiting solution and its derivatives up to second order, which allows them to extend the strategy given in [24], to $\mathbb{R}^{d}$.
@@
Quantitative estimates in the whole space were derived, in the context of moderately interacting particles  [^2] The moderately interacting particle model was introduced by Oelschläger in [35, 34, 36]. with singular kernels and without common noises - including the Biot-Savart, the Keller-Segel and the Dirac delta measure kernels - by [37] and [38], based on a semigroup approach introduced in [18], with further advancements in [8], [22], [28] and [39]. It is important to note that these works do not rely on the relative entropy method, but instead establish quantitative estimates in strong functional topologies. The connection between the time evolution of the relative entropy, of the joint distribution, and moderate interacting regime, on the whole space was carried out by [10], based on a combination between the relative entropy and the regularised $L^{2}$-estimate in [36], deriving a propagation of chaos result for the viscous porous medium equation.
@@
However, it was in [40] that equation (1) was derived on the torus for singular kernels and common noise in the moderately interacting regime, by applying the relative entropy directly to the mollified empirical measure. Earlier, also in the same domain, the results established in [37] were extended to the setting with environmental noise in [27]. It is worth highlighting, that quantitative entropy estimates for systems with individual and common noise, have been obtained, as shown in [44] for incompressible the Navier-Stokes equations, [9] for the Hegselmann-Krause model, and [33] for mean-field systems with bounded kernels. For more results, see [19], [15], [16], and [29]. Particle systems driven by environmental noises, such as given by (2), appear in several contexts, such as problems in statistical mechanics, population dynamics and biology, as well as in the analysis of numerical Monte–Carlo schemes and related applications; see, e.g., [4], [5], [25], [20], and [30].
@@
Outline of the article
In order to present our results, this work is organized as follows. In the next two subsections, 1.1 and 1.2, respectively, we introduce the notations and assumptions used throughout the text, setting the stage for the subsequent discussions. In the subsection 1.3, we present our main results. The sections 2 and 3 are dedicated to the proofs of our main results. To ease the presentation, the Appendix, provides the existence of a suitable solution for the limiting equation, as well as some inequalities used throughout the proofs.
@@ 1.1 Notations
For $d\geq 1$, the space of probability density functions on $\mathbb{R}^{d}$ is denoted by $\mathcal{P}(\mathbb{R}^{d})$. For a normed vector space $U$, we denote its norm by $\|\cdot\|_{U}$, except for $L^{p}$ spaces, where we write $\|\cdot\|_{p}$, for $p\geq 1$.
For $\gamma\in(0,1]$ the Hölder space on $\mathbb{R}^{d}$ is given by
$C^{\gamma}=C^{\gamma}(\mathbb{R}^{d})\doteq\left\{f:\mathbb{R}^{d}\to\mathbb{R}^{e}\mid\|f\|_{\gamma}\doteq\|f\|_{\infty}+\sup_{x,y\in\mathbb{R}^{d}}\frac{|f(x)-f(y)|}{|x-y|^{\gamma}}<\infty\right\}.$
The Kantorovich-Rubinstein metric reads, for any two probability measures $\mu$ and $\nu$ on $\mathbb{R}^{d}$,
@@ 1. $(\mathbf{A}^{V})$
    Let $V^{N}:\mathbb{R}^{d}\to\mathbb{R}$ be the scaling given by $V^{N}(y)=N^{\beta}V(N^{\frac{\beta}{d}}y)$, $\beta\in(0,1)$, with $V\in\mathcal{P}\left(\mathbb{R}^{d}\right)$ satisfying
@@
    for $\alpha>d/2$
    $V(y)\leq C_{d}(1+|y|^{2})^{-\alpha},$ (6)
    and for $\gamma\in(0,1)$ and $y\neq 0$
    $V(y)\leq C_{d}|y|^{-3d-\gamma}.$ (7)
  2. $(\mathbf{A}^{K})$
    It holds for $\gamma\in(1/2,1)$ that
    $\|K\ast f\|_{\gamma}\leq C\left(\|f\|_{1}+\|f\|_{4}\right)$ (8)
    for all $f\in H^{1}$.
@@ 5. $(\mathbf{A}^{\rho_{0}})$
    The initial condition $\rho_{0}$ is taken such that $\rho_{0}\in W^{2,\infty}\cap\mathcal{P}\left(\mathbb{R}^{d}\right)$ and
@@ 6. $(\mathbf{A}^{X_{0}})$
    The initial particle system $(X_{0}^{i,N})_{i=1}^{N}$, $N\in\mathbb{N}$ is a sequence of i.i.d random variables with law $\rho_{0}$.
@@
For an example of mollifier $V\in\mathcal{P}(\mathbb{R}^{d})$ satisfying $\left(\mathbf{A}^{V}\right)$, take
@@
and then, it is the divergence in the sense of distributions, of an $L^{\infty}$ matrix $K_{0}$, namely $K=\nabla\cdot K_{0}$, with
$K_{0}(x_{1},x_{2})=-\left(\frac{1}{2\pi}\arctan{\left(\frac{x_{1}}{x_{2}}\right)}\right)I.$ (14)
It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$. Furthermore, $K$ satisfy $(\mathbf{A}^{K})$, see [39].
@@ 1.3 Statement of the main results
In this subsection, we establish our main results. We begin by formulating a suitable notion of solution, needed for deriving quantitative entropy estimates.
Definition 1.1.
Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$. A pathwise solution for (1) with initial condition $\rho_{0}$, is a measurable function $\rho:[0,T_{1}]\times\mathbb{R}^{d}\times\Omega\to\mathbb{R}$, for some $T_{1}>0$, such that:
  • it holds $\mathbb{P}$-a.s.
    $\rho\in L^{\infty}\left([0,T_{1}];W^{2,1}\cap W^{2,\infty}\cap\mathcal{P}(\mathbb{R}^{d})\right)$ (15)
@@
    for all $t\in[0,T_{1}]$ and constants $C_{1},C_{2}$ and $C_{3}$ dependent on $\widetilde{C}_{1}$, $\widetilde{C}_{2}$ and $\widetilde{C}_{3}$.
@@
Assume $(\mathbf{A}^{V})$, $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, $(\mathbf{A}^{\rho_{0}})$, $(\mathbf{A}^{X_{0}})$ and let $\rho$ be a solution of (1), given by Definition 1.1. In addition, let the dynamics of the particle system be given by (2) and
$\lim_{N\to\infty}N^{\theta}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.$ (20)
and
@@
Then, there exist a time $T\in(0,T_{1})$, an almost surely finite random variable $A_{0}$, and a set $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the following property: for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (21)
In particular, we get
$\limsup_{N\to\infty}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim T,\,\,\,\,\,\,\mathbb{P}-a.s.$ (22)
The following result establishes energy estimates as an application of our entropy bounds, extending previous results by [27] and [37]. Precisely, we utilize the entropy dissipation to control the Fisher information of the particle system, which in turn allows us to derive, a new energy estimate for the difference between the regularized empirical measure and the solution to (1), using the Ladyzhenskaya inequality and a nonlinear Gronwall lemma.
Theorem 2.
Assume $(\mathbf{A}^{V})$, $(\mathbf{A}^{K})$, $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, $(\mathbf{A}^{\rho_{0}})$, $(\mathbf{A}^{X_{0}})$, $d=2$ and let $\rho$ be a solution of (1), given by Definition 1.1. In addition, let the dynamics of the particle system be given by (2) and
$\lim_{N\to\infty}N^{\widetilde{\theta}}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.,$ (23)
$\lim_{N\to\infty}N^{\widetilde{\theta}}\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}=0,\,\,\,\mathbb{P}-a.s.$ (24)
and
$\widetilde{\theta}\doteq\min{\left(\theta_{1},\theta_{2},\widetilde{\theta}_{1},\widetilde{\theta}_{2},\widetilde{\theta}_{3},\widetilde{\theta}_{4}\right)}-\delta$
where
$\theta_{1}\doteq\left(1-\beta\left(2+2\alpha\right)\right);\,\,\,\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{2}\Big)\right);\,\,\,\widetilde{\theta}_{1}\doteq\left(1-2\beta(1+\frac{1}{2})+\beta\right)$
$\widetilde{\theta}_{2}\doteq\beta(2-3\gamma);\,\,\,\widetilde{\theta}_{3}\doteq\frac{\beta\gamma^{2}}{2};\,\,\,\widetilde{\theta}_{4}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{2}\Big)\right)$
with $\gamma\in(1/2,2/3)$ and $\delta>0$, such that $\widetilde{\theta}>0$ and $\beta\in\left(0,\frac{1}{2+2\alpha}\right)$.
Then, there exist a time $T\in(0,T_{1})$, a pair of almost surely finite random variables $A_{0}$, $\widetilde{A}_{0}$, and a set $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the following property: for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have
$\sup_{t\in[0,T]}\|\rho_{t}^{N}-\rho_{t}\|_{2}^{2}+\int_{0}^{T}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
[…]
In particular, we get
$\limsup_{N\to\infty}\left(\sup_{t\in[0,T]}\|\rho_{t}^{N}-\rho_{t}\|_{2}^{2}+\int_{0}^{T}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds\right)$ $\lesssim T,\,\,\,\mathbb{P}-a.s.$ (26)
As a corollary of Theorem 1, we establish quantitative bounds for the marginals of the empirical measure of the particle system. In the sense of [1, Section 8.3], it corresponds to get bounds of the empirical measure in the sense to the Kantorovich–Rubinstein metric, given in (4).
@@
Under the assumptions of Theorem 1, for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$, such that for $N\geq N_{0}$ we have
$\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (27)
In particular, we get
$\limsup_{N\to\infty}\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim T,\,\,\,\mathbb{P}-a.s.$ (28)
@@
We observe that, compared with previous works on moderate particle systems without common noise, we lose the order of convergence. This is because the estimation of the martingale term in our work is is of order $\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)$, while in [37] is the order $\frac{1}{2}-\frac{\beta}{2}$, see Theorem 1.3 in [37]. When environmental noise is taken into account, the overall framework remains comparable to that of [27]. For the same reason, the range of the parameter $\beta$ is smaller in our work. However, in the Theorem 1, we obtain an improved convergence rate compared to that established in [40].
The next result, whose proof is postponed to the Appendix, establishes that the limiting equation is well posed in the appropriate functional framework, given by Definition 1.1.
Theorem 3.
Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$. Then, there exists a solution of (1) in the sense of Definition 1.1.
2 Proofs of Theorem 1
In this section we present the proofs of Theorem 1. In Subsection 2.1, by applying the Itô’s formula, such as in [40], we derive an evolution equation for the relative entropy functional of the regularized empirical measure with respect to the solution of (1). In Subsection 2.2, we deal with the quadratic variation terms arising from the application of Itô’s formula. In Subsection 2.3, we address the nonlinear contributions appearing in the evolution equation for the relative entropy functional obtained in the first step.